OpenStudy (anonymous):

What is the slope of the line 5x - 2y = 10?

OpenStudy (phi):

you want to change this to y = mx+b m (the number in front of the x) will be the slope first step: add -5x to both sides, what do you get?

OpenStudy (anonymous):

-2y = 10 + 5x right?

OpenStudy (phi):

-5x to both sides try again: -5x + 5x - 2y = 10 -5x simplify it

OpenStudy (phi):

now divide both sides by -2 (that means every term on both sides get divided by -2) what do you get?

OpenStudy (anonymous):

1st make y the subject of the formula... 5x-2y=10 -2y=10-5x y=(10-5x)/2 Or y= (10/2) - (5/2)x therefor, y=mx+c has m as gradient (slope) so, 5/2 Or 2.5 is the Slope